5. A Ping-Pong ball has a diameter of 3.80 cm and average density



of 0.084 0 g/cm3. What force is required to hold it completely



submerged under water?

1
Expert's answer
2022-06-13T08:24:09-0400

Given:

"\rho=0.0840\:\rm g/cm^3"

"\rho_w=1.00\:\rm g/cm^3"

"R=1.90\:\rm cm"


"F=F_b-mg=g\rho_w V-mg=gV(\rho_w-\rho)"

"F=9.8*\frac{4}{3}*3.14*1.90^3*(1.00-0.0840)*\frac{1}{1000}\\
=0.258\:\rm N"